LORENA BOUGHT 10 PACKS OF STYROFOAM CUPS FOR THE GRADUATION DANCE.
A PACK OF FIFTY 12-0Z CUPS COSTS $1.80 AND A PACK OF FIFTY 16OZ CUPS COSTS $2.40.
LORENA PAID A TOTAL OF $21.60 EXCLUDING TAX.
HOW MANY PACKS OF EACH SIZE CUP DID SHE BUY?

Answer by checkley77(12844)

2.40X+1.80(10-X)=21.60
2.40X+18-1.8X=21.60
.6X=21.60-18
.6X=3.6
X=3.6/.6
X=6 PACKS OF $2.40 CUPS.
10-6=4 PACKS OF $1.80 CUPS.
PROOF;
2.40*6+1.8*4=21.6
14.4+7.2=21.6
21.6=21.6